OpenStudy (anonymous):

A bag contains 2 blue and 3 green marbles. Two marbles are drawn at random with replacement. Which of the following gives the probability of drawing two blue marbles?

OpenStudy (nikolas):

Try drawing out a tree diagram for these kinds of problems if it helps. For the first draw there are 2 blue marbles, and 5 marbles in total, so the probability that you will draw a blue marble is 2/5. For the next draw, if you already removed one blue marble, there is 1 blue marble remaining and 4 blue marbles in total, so the probability that you draw another blue marble is 1/4. You can multiply these two probabilities together to get the total probability of drawing two: 1/5 * 2/4 = 2/20 = 1/10. This exact problem is described, along with more helpful info, here: http://www.mathsisfun.com/data/probability-events-conditional.html